# arXiv:2603.01209v1[cs.AI] 01 Mar 2026

### v1

@@ Abstract
Tool-augmented LLM agents increasingly solve tasks by interleaving natural-language reasoning with executable Python actions. Many agent frameworks equip models with a persistent interpreter, but training traces often leave this execution contract implicit. We ask whether interpreter persistence is merely a runtime scaffold, or a training-time semantic that shapes how agents learn to use the interpreter.
We answer with a controlled $2\times 2$ study, conducted on a single task family and base model, crossing training semantics (fine-tuning on persistent vs. stateless-execution traces) and runtime semantics (evaluating in a persistent vs. stateless interpreter). We introduce Opaque Knapsack, a deliberately non-collapsible, partially observable knapsack variant with budgeted tool access. For each instance, we generate matched trajectories differing only in the persistence/reset execution contract, fine-tune identical models on each variant, and evaluate all four train-runtime combinations.
We find that misaligning training and runtime semantics produces two characteristic failure modes: a model trained on persistent traces but deployed in a stateless interpreter triggers missing-variable errors in ${\sim}80\%$ of episodes, entering cascading recovery loops that consume its token budget without making progress. The reverse—a stateless-trained model deployed in a persistent runtime—pays an amnesia tax: it redundantly re-derives state that the interpreter could have retained, using roughly $3.5\times$ more tokens than the aligned persistent condition.
Notably, these efficiency and stability costs are not accompanied by a statistically significant reduction in solution quality—execution semantics primarily affect how agents reach solutions, not whether they do. This suggests that aligning fine-tuning traces with the deployment runtime reduces inference footprint without sacrificing task performance, and that interpreter persistence should be treated as a first-class semantic of agent datasets rather than a hidden implementation detail at runtime.
@@ 1 Introduction
Tool-augmented language models are increasingly deployed as agents that solve tasks by interleaving natural-language deliberation with executable actions—often Python code executed in an interpreter Wang et al. (2024b); Roucher et al. (2025); Gao et al. (2023)—and then reacting to environment feedback Yao et al. (2023). In these systems, the interpreter is not just a calculator: it is a workspace where the agent can accumulate variables, data structures, and intermediate results across turns.
This execution contract varies across agent frameworks. CodeAct-style agents typically run inside a persistent Python session Wang et al. (2024b), while ReAct-style agents externalize state into the context window Yao et al. (2023), and other systems maintain dedicated memory stores Packer et al. (2024). Yet training pipelines usually treat agent traces as token sequences and leave execution semantics implicit. This mismatch motivates our research question:
Is interpreter persistence merely an inference-time scaffold, or a training-time semantic that shapes how agents learn to use tools?
The distinction matters in practice: Agent builders routinely fine-tune models on tool-use logs produced by a particular runtime and then deploy the resulting model in a (sometimes different) runtime. If execution semantics are learnable, then matching training traces to deployment semantics can determine whether a CodeAct agent exploits persistent state and how efficiently it does so under latency and cost constraints.
We answer this question with a controlled $2\times 2$ study that disentangles training semantics from runtime scaffolding, conducted on a single task family and base model. We introduce Opaque Knapsack, a partially observable knapsack variant designed to be non-collapsible for executable-action agents: Item attributes and feasibility constraints are hidden behind budgeted tool calls, and per-step interaction limits force iterative information gathering and state revision. For each task instance, we generate paired interleaved reasoning-action trajectories matched by task instance, prompt, tool interface, and supervision, differing only in the persistence/reset execution contract: whether the interpreter state persists across turns or is reset after each action. We fine-tune identical Qwen3-8B models on each trace type and evaluate them under both runtime semantics.
@@
  • Execution semantics as a training variable. We identify interpreter persistence as part of the agent trace contract, studying execution semantics as learnable components of the training signal with a controlled $2\times 2$ train/runtime design.
@@
  • Evidence that persistence is learned. Using a $2\times 2$ train/runtime cross-evaluation, we show that exposure to persistence during fine-tuning shapes state-management behavior. These results suggest that interpreter persistence should be treated as a first-class semantic of agent datasets and training protocols, not a hidden implementation detail of the runtime. Relatedly, aligning the execution semantics of training traces with the deployment runtime lowers inference cost without compromising performance.
@@ 2 Related Work
A growing body of work studies agents that interleave natural-language reasoning with executable actions. Program-aided and executable-action frameworks such as PAL (Gao et al., 2023), PoT (Chen et al., 2023b), ToRA (Gou et al., 2024), and CodeAct (Wang et al., 2024b; Roucher et al., 2025) demonstrate that offloading computation to an external interpreter can improve reliability on multi-step tasks. Interleaved reasoning–action paradigms such as ReAct (Yao et al., 2023) further formalize agents as systems that alternate between deliberation and interaction. Recently, the Agent–Computer Interface (ACI) perspective has emphasized that the interface between an agent and its execution environment—including tool affordances, feedback formats, and an interactive runtime—is itself a key design surface for agent performance. For example, SWE-agent (Yang et al., 2024) shows that purpose-built ACIs for software engineering tasks substantially improve agent effectiveness. However, ACI work largely treats the execution environment as an inference-time systems choice. In contrast, we isolate a specific execution semantic—interpreter state persistence—and study it as part of the training-time trace contract, measuring how training/runtime alignment shapes learned state-management behavior.
@@
A recurring theme in agent learning is that capabilities needed at deployment must appear in the training distribution. Toolformer (Schick et al., 2023) shows that effective tool invocation emerges when tool use is embedded in training, and recent work argues that the allocation and timing of reasoning should align between training and inference (Nguyen et al., 2026; Akter et al., 2025). We extend this alignment principle to execution semantics: our results suggest that if persistent state is available at deployment, exposing the same persistence contract during fine-tuning can shape whether (and how) agents learn to delegate state to the interpreter. Conversely, mismatching training trace semantics and deployment runtimes can induce compounding behavioral distribution shift, a known vulnerability in imitation learning (Ross et al., 2011).
@@ 4.1.1 Datasets and Trace Generation
Few-shot demonstrations and contract consistency.
Each runtime regime includes a small number of format-only demonstrations whose sole purpose is to make the execution contract concrete (persistent reuse vs. mandatory redefinition / explicit printing of state). These demonstrations are intentionally task-agnostic (they do not contain knapsack-specific information) and are treated as part of the runtime interface: using persistence-style examples in a stateless interpreter (or vice versa) would create an internally inconsistent protocol. At evaluation time the prompt—including the demonstrations and runtime instruction—is held fixed within each runtime. Thus, differences we report between persistent-trained and stateless-trained models under the same runtime (e.g., token footprint under a persistent interpreter) cannot be attributed to different demonstrations at inference. Moreover, in the mismatch condition the persistent-trained model is explicitly shown stateless-style demonstrations yet still produces missing-binding errors characteristic of reliance on cross-turn executable state, indicating that the dependence is learned rather than merely prompted.
@@ 5 Discussion
Conversely, persistent execution makes it advantageous to keep task-relevant state as executable bindings with non-zero lifespan in the interpreter, reducing redundancy and enabling shorter, more stable interaction traces. Importantly, the stateless regime is not data-limited (as shown in Table 1): It produces longer, more tool-heavy trajectories and substantially more tokens during trace generation, yet yields a substantially larger interaction footprint and directionally lower solution quality downstream. This indicates that token volume is an unreliable proxy for learnability in agent datasets; the execution contract can make traces qualitatively easier or harder to learn from even when tasks, tools, and supervision are held fixed. This echoes findings that SFT teaches behavioral style and format as much as task capability (Zhou et al., 2023) — here, the execution contract is the stylistic prior being absorbed.
@@
Misalignment between what the model was trained to assume and what the runtime actually provides induces characteristic failures. When a persistent-trained model is evaluated under a stateless runtime, it produces missing-binding exceptions in roughly $80\%$ of episodes and frequently enters error-dense recovery loops (Section 4.3). These failures are asymmetric: They do not appear in the other train/runtime configurations, indicating that persistent-trained policies are not merely “helped” by statefulness but depend on it as part of their learned execution strategy. From an engineering perspective, this suggests a concrete risk: swapping runtimes (or silently changing interpreter persistence) can degrade both correctness and stability without any change to the model weights.
Implications for dataset and system design.
Two implications follow: First, execution semantics (e.g., persistence vs. statelessness as well as how state is exposed or summarized) should be treated as first-class attributes of agent datasets and training protocols, alongside tool schemata and supervision signals. Second, benchmarks for executable-action agents should explicitly prevent collapse into one-shot scripts; Opaque Knapsack provides one such blueprint by combining tool-gated hidden attributes, hard interaction budgets, and non-monotonic constraints that force iterative revision.
Falsifiability.
If interpreter persistence were only an inference-time scaffold, we would expect little or no interaction between training semantics and runtime semantics in our design. Specifically, a persistent runtime should reduce token cost similarly regardless of which adapter is deployed; a persistent-trained model should transfer cleanly to a stateless interpreter when shown the stateless contract, without systematic missing-binding errors; and the base model should benefit materially from persistent execution once tool access is available. Any of these outcomes would weaken the conclusion that persistence is a learned behavioral prior that must be aligned between training traces and the deployment runtime. Instead, we observe the opposite in each case: Token efficiency gains are concentrated in the aligned persistent condition, missing-binding errors are asymmetric and specific to the persistent-trained/stateless-runtime mismatch, and the base model remains near-floor regardless of runtime statefulness—which is consistent with persistence functioning as a training-time semantic rather than a post-hoc scaffold.
@@ Appendix B Agent Implementation
To generate valid trajectories interleaving reasoning and executable actions, we implemented a custom CodeAct-style agent loop. This appendix documents the execution contract, runtime alignment across regimes, and the exact system prompts used in the experiments.
@@ B.5.2 Runtime-State Instructions
To explicitly define the execution contract for the model, the base system prompt is appended with a specific runtime instruction that varies by regime. Listing 3 details the instruction for the persistent environment, while Listing 4 outlines the corresponding reset instruction for the stateless environment.
@@ B.5.3 Few-Shot Demonstrations
In addition to the base system prompt and runtime-state instruction, the agent is provided with a small number of format demonstrations at the beginning of each episode. These demonstrations illustrate the required Reflection–Action–Observation structure and the handling of runtime state. They are intentionally task-agnostic and do not introduce any Opaque Knapsack-specific information; their sole purpose is to make the execution contract concrete. The few-shot content differs between regimes to faithfully reflect the underlying interpreter semantics. Representative demonstrations for each regime are shown below; additional examples follow the same pattern.
@@
The demonstrations are identical in task complexity and formatting across regimes and do not introduce additional task information. They serve only to illustrate the execution contract enforced by the runtime. The interpreter independently enforces persistence or stateless behavior regardless of the prompt; the demonstrations do not alter the underlying execution semantics. In the mismatch condition, the persistent-trained model is explicitly shown stateless-style demonstrations at evaluation time yet still produces missing-binding errors characteristic of cross-turn state reliance, confirming that the learned dependence is behavioral rather than merely prompted.

### v2

# Agents Learn Their Runtime: Interpreter Persistence as Training-Time Semantics

@@ Abstract
Tool-augmented LLM agents increasingly solve tasks by interleaving natural-language reasoning with executable Python actions. Many agent frameworks equip models with a persistent interpreter, but training traces often leave this runtime assumption implicit. We ask whether interpreter persistence is merely a runtime scaffold, or a property of the training data that shapes how agents learn to use the interpreter.
We answer with a controlled $2\times 2$ study, conducted on a single task family and base model, crossing training conditions (fine-tuning on persistent vs. stateless-execution traces) and runtime conditions (evaluating in a persistent vs. stateless interpreter). We introduce Opaque Knapsack, a deliberately non-collapsible, partially observable knapsack variant with budgeted tool access. For each instance, we generate matched trajectories differing only in whether the interpreter state persists or resets between turns, fine-tune identical models on each variant, and evaluate all four train-runtime combinations.
We find that misaligning training and runtime conditions produces two characteristic failure modes: a model trained on persistent traces but deployed in a stateless interpreter triggers missing-variable errors in ${\sim}80\%$ of episodes, entering cascading recovery loops that consume its token budget without making progress. The reverse—a stateless-trained model deployed in a persistent runtime—pays an amnesia tax: it redundantly re-derives state that the interpreter could have retained, using roughly $3.5\times$ more tokens than the aligned persistent condition.
Notably, these efficiency and stability costs are not accompanied by a statistically significant reduction in solution quality–interpreter persistence shapes how agents reach solutions, not whether they do. This suggests that the runtime used to generate fine-tuning traces should be an explicit design choice, not a hidden implementation detail.
@@ 1 Introduction
Tool-augmented language models are increasingly deployed as agents that solve tasks by interleaving natural-language deliberation with executable actions—often Python code executed in an interpreter Wang et al. (2024b); Roucher et al. (2025); Gao et al. (2023)—and then reacting to environment feedback Yao et al. (2023). In these systems, the interpreter is a workspace where the agent can accumulate variables, data structures, and intermediate results across turns.
Agent frameworks differ in how they manage interpreter state. CodeAct-style agents Wang et al. (2024b) run inside a persistent Python session where variables accumulate across turns, but the traces used to post-train the underlying models rarely make this explicit. This mismatch motivates our research question:
Is interpreter persistence merely an inference-time scaffold, or a property of training that shapes how agents learn to use tools?
The distinction matters in practice: models are routinely fine-tuned on traces generated in one runtime and then deployed in another. Whether the model absorbed the runtime’s behavior during post-training determines how efficiently and how reliably it operates at deployment.
We answer this question with a controlled $2\times 2$ study, conducted on a single task family and base model, that disentangles what the model learned during post-training from what the runtime provides at deployment. We introduce Opaque Knapsack, a partially observable knapsack variant designed to be non-collapsible for executable-action agents: Item attributes and feasibility constraints are hidden behind budgeted tool calls, and per-step interaction limits force iterative information gathering and state revision. For each task instance, we generate paired interleaved reasoning-action trajectories matched by task instance, prompt, tool interface, and supervision, differing only in whether the interpreter state persists or resets between turns: whether the interpreter state persists across turns or is reset after each action. We fine-tune identical Qwen3-8B models on each trace type and evaluate them under both runtime semantics.
@@
  • Evidence that persistence is learned. Using a $2\times 2$ train/runtime cross-evaluation, we show that interpreter persistence is absorbed as a behavioral prior during post-training — carrying into deployment regardless of prompting, and producing characteristic failure modes when training and runtime are misaligned.
@@ 2 Related Work
A growing body of work studies agents that interleave natural-language reasoning with executable actions. Program-aided and executable-action frameworks such as PAL (Gao et al., 2023), PoT (Chen et al., 2023b), ToRA (Gou et al., 2024), and CodeAct (Wang et al., 2024b; Roucher et al., 2025) demonstrate that offloading computation to an external interpreter can improve reliability on multi-step tasks. Interleaved reasoning–action paradigms such as ReAct (Yao et al., 2023) further formalize agents as systems that alternate between deliberation and interaction. Recently, the Agent–Computer Interface (ACI) perspective has emphasized that the interface between an agent and its execution environment—including tool affordances, feedback formats, and an interactive runtime—is itself a key design surface for agent performance. For example, SWE-agent (Yang et al., 2024) shows that purpose-built ACIs for software engineering tasks substantially improve agent effectiveness. However, ACI work largely treats the execution environment as an inference-time systems choice. In contrast, we isolate a specific execution semantic—interpreter state persistence—and study it as part of the training-time trace semantics, measuring how training/runtime alignment shapes learned state-management behavior.
@@
A recurring theme in agent learning is that capabilities needed at deployment must appear in the training distribution. Toolformer (Schick et al., 2023) shows that effective tool invocation emerges when tool use is embedded in training, and recent work argues that the allocation and timing of reasoning should align between training and inference (Nguyen et al., 2026; Akter et al., 2025). We extend this alignment principle to execution semantics: our results suggest that if persistent state is available at deployment, exposing the same persistence behavior during fine-tuning can shape whether (and how) agents learn to delegate state to the interpreter. Conversely, mismatching training trace semantics and deployment runtimes can induce compounding behavioral distribution shift, a known vulnerability in imitation learning (Ross et al., 2011).
@@ 4.1.1 Datasets and Trace Generation
Few-shot demonstrations and semantics consistency.
Each runtime regime includes a small number of format-only demonstrations whose sole purpose is to make the execution semantics concrete (persistent reuse vs. mandatory redefinition / explicit printing of state). These demonstrations are intentionally task-agnostic (they do not contain knapsack-specific information) and are treated as part of the runtime interface: using persistence-style examples in a stateless interpreter (or vice versa) would create an internally inconsistent protocol. At evaluation time the prompt—including the demonstrations and runtime instruction—is held fixed within each runtime. Thus, differences we report between persistent-trained and stateless-trained models under the same runtime (e.g., token footprint under a persistent interpreter) cannot be attributed to different demonstrations at inference. Moreover, in the mismatch condition the persistent-trained model is explicitly shown stateless-style demonstrations yet still produces missing-binding errors characteristic of reliance on cross-turn executable state, indicating that the dependence is learned rather than merely prompted.
@@ 5 Discussion
Conversely, persistent execution makes it advantageous to keep task-relevant state as executable bindings with non-zero lifespan in the interpreter, reducing redundancy and enabling shorter, more stable interaction traces. Importantly, the stateless regime is not data-limited (as shown in Table 1): It produces longer, more tool-heavy trajectories and substantially more tokens during trace generation, yet yields a substantially larger interaction footprint and directionally lower solution quality downstream. This indicates that token volume is an unreliable proxy for learnability in agent datasets; the execution semantics can make traces qualitatively easier or harder to learn from even when tasks, tools, and supervision are held fixed. This echoes findings that SFT teaches behavioral style and format as much as task capability (Zhou et al., 2023) — here, the execution semantics is the stylistic prior being absorbed.
@@
Misalignment between what the model was trained to assume and what the runtime actually provides induces characteristic failures. When a persistent-trained model is evaluated under a stateless runtime, it produces missing-binding exceptions in roughly $80\%$ of episodes and frequently enters error-dense recovery loops (Section 4.3). These failures are asymmetric: They do not appear in the other train/runtime configurations, indicating that persistent-trained policies are not merely “helped” by statefulness but depend on it as part of their learned execution strategy. This suggests a concrete risk: swapping runtimes (or silently changing interpreter persistence) can degrade both correctness and stability without any change to the model weights.
Ruling out the scaffold interpretation.
These results contradict the view that persistence is just a runtime scaffold. If it were, a persistent runtime should reduce token cost regardless of how the model was trained, and a persistent-trained model should adapt cleanly when shown stateless demonstrations at inference. Neither holds: efficiency gains are concentrated in the aligned persistent condition while missing-binding errors appear only in the persistent-trained/stateless-runtime mismatch.
@@ Appendix B Agent Implementation
To generate valid trajectories interleaving reasoning and executable actions, we implemented a custom CodeAct-style agent loop. This appendix documents the execution semantics, runtime alignment across regimes, and the exact system prompts used in the experiments.
@@ B.5.2 Runtime-State Instructions
To explicitly define the execution semantics for the model, the base system prompt is appended with a specific runtime instruction that varies by regime. Listing 3 details the instruction for the persistent environment, while Listing 4 outlines the corresponding reset instruction for the stateless environment.
@@ B.5.3 Few-Shot Demonstrations
In addition to the base system prompt and runtime-state instruction, the agent is provided with a small number of format demonstrations at the beginning of each episode. These demonstrations illustrate the required Reflection–Action–Observation structure and the handling of runtime state. They are intentionally task-agnostic and do not introduce any Opaque Knapsack-specific information; their sole purpose is to make the execution semantics concrete. The few-shot content differs between regimes to faithfully reflect the underlying interpreter semantics. Representative demonstrations for each regime are shown below; additional examples follow the same pattern.
@@
The demonstrations are identical in task complexity and formatting across regimes and do not introduce additional task information. They serve only to illustrate the execution semantics enforced by the runtime. The interpreter independently enforces persistence or stateless behavior regardless of the prompt; the demonstrations do not alter the underlying execution semantics. In the mismatch condition, the persistent-trained model is explicitly shown stateless-style demonstrations at evaluation time yet still produces missing-binding errors characteristic of cross-turn state reliance, confirming that the learned dependence is behavioral rather than merely prompted.